OpenStudy (anonymous):

ANSWER how many basketball games are played in a 10-team league if each team plays all other team TWICE

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

180?

OpenStudy (anonymous):

someone should confirm that for me

OpenStudy (anonymous):

ANSWER how many basketball games are played in a 10-team league if each team plays all other team TWICE

OpenStudy (anonymous):

@AravindG, are you trying to scare this person off? Maybe they are just having problems with some homework.

OpenStudy (anonymous):

PLZ

OpenStudy (aravindg):

Nope I was welcoming the user and also making sure that he.she gets to know how the site works . The non cheating policy is an important feature of our site @kashbmaryd

OpenStudy (anonymous):

Deenal answer

OpenStudy (anonymous):

ANSWER how many basketball games are played in a 10-team league if each team plays all other team TWICE

OpenStudy (anonymous):

Understood. Just seems a little pushy to hit him/her with the non-cheating policy on the first question. But anyway, I still like the site @ AravindG

OpenStudy (anonymous):

it should be 90!

OpenStudy (anonymous):

\[10C2*2\]

OpenStudy (kropot72):

The number of combinations of the 10 teams is 10C2 which is\[\frac{10!}{2!8!}=\frac{10\times 9}{2}=45\] Therefore there would be 45 games in the first round.

OpenStudy (anonymous):

it should be 90 games!!

OpenStudy (kropot72):

And another 45 games in the second round.

OpenStudy (anonymous):

what is the answer @JOJP ?

OpenStudy (kropot72):

making 45 + 45 = 90 games.

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

see choosing any 2 teams to play a match from the 10 teams is 10C2 OK!! now each team is playing 2 mathces so we have 10C2*2

OpenStudy (anonymous):

which is equal to 90

OpenStudy (anonymous):

ok thanks alooot all

OpenStudy (anonymous):

Seven seats are positioned in a row at a movie theater. Rebecca, Wayne, Heather, Garrett, Mark, Ashley, and Alice want to sit together in this row. a) How many different ways can they be arranged? b)How many different ways can they be arranged if Rebecca sits in the second seat? c) How many different ways can they be arranged if Garrett sits in the left aisle seat and Mark sits in the right aisle seat?

OpenStudy (anonymous):